OpenStudy (anonymous):

Which of the following are solutions to the equation x^2 - 49 = 0? *Check all that apply* A. -7 B. 2 C. -14 D. 49 E. 7 F. 14

OpenStudy (anonymous):

x^2 - 49 = (X+7)(X-7)

OpenStudy (anonymous):

So it's 7 and -7?

OpenStudy (jdoe0001):

keep in mind that \(\bf (a^2-b^2) = (a-b)(a+b)\)

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

(X+7)(X-7) = 0 x+7 = 0 => x = -7 x-7 = 0 => x = 7

OpenStudy (anonymous):

Just 7 and -7? I hate it when they give me 18,000 other options -.- smh

OpenStudy (jdoe0001):

well, the idea is not to pick one out, the idea is that you evaluate the expression, then it doesn't matter how many they give you, you know which is which

OpenStudy (anonymous):

True... lol much apprectiation though thanks